OpenStudy (anonymous):

18 is what percent of 50?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

A. 3600% B. 0.36% C. 36% D. 3.6%

OpenStudy (anonymous):

C.36%

OpenStudy (anonymous):

oh 18 is what percent of 50 sorry

OpenStudy (anonymous):

\[\frac{18}{50}=\frac{36}{100}=36\%\]

OpenStudy (anonymous):

36%

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

proportions. 50 = 1/2 of 100 and 18 = 1/2 of 36

OpenStudy (anonymous):

ok, so you would have to do percent over 100 what they did is do 18/50 but we don't know the percent, so we would put an x over 100 now we got 18/50 = x/100 then we would solve like a proportion 18 * 100 = 1800 then 50 * x =50x now we got 50x = 1800 we can divide by 50 to solve for x we get x = 36

OpenStudy (anonymous):

Wonderful thank you so much !